### v1

# Towards General Conceptual Model Editing via Adversarial Representation Engineering

Abstract
Recent research has introduced Representation Engineering (RepE) as a promising approach for understanding complex inner workings of large-scale models like Large Language Models (LLMs). However, finding practical and efficient methods to apply these representations for general and flexible model editing remains an open problem. Inspired by the Generative Adversarial Network (GAN) framework, we introduce a novel approach called Adversarial Representation Engineering (ARE). This method leverages RepE by using a representation sensor to guide the editing of LLMs, offering a unified and interpretable framework for conceptual model editing without degrading baseline performance. Our experiments on multiple conceptual editing confirm ARE’s effectiveness. Code and data are available at https://github.com/Zhang-Yihao/Adversarial-Representation-Engineering.
1 Introduction
Figure 1: An illustration of ARE framework. This example showcases how ARE can enhance the concept of "angry" within an LLM. The process involves an iterative dance between the generator and the discriminator. The generator produces outputs, while the discriminator refines its internal representation of "angry" based on these outputs. Through this back-and-forth training, the LLM gradually learns to produce outputs that align better with the concept of "angry."
While Large Language Models (LLMs) have achieved remarkable success in a variety of tasks [37], their complex internal mechanism makes interpreting and censoring their behaviors (e.g., for safety alignment or hallucination reduction) challenging.
To improve the interpretability and consequently the safety of LLMs, numerous efforts have been dedicated to interpreting the internal mechanisms from various perspectives like feature attribution [11, 44], neuron analysis [16, 35], and self-explanation [20].
Recently, Zou et al. proposed the idea of Representation Engineering (RepE[71], which offers a way of understanding how LLMs work internally by focusing on the overall feature representations rather than individual neurons. Specifically, RepE extracts and analyzes the intermediate features of different concepts (e.g., honesty, fairness, and harmlessness), enabling the monitoring of the internal behaviors of LLMs. More relevantly, RepE potentially allows editing and controlling the behaviors of LLMs by directly intervening in the internal hidden layers during inference.
However, as RepE was essentially proposed to monitor the behaviors of LLMs, their proposed method for editing the model through representation vector incorporation is rather limited for practical uses. For instance, their method could disrupt the underlying structure of general LLMs, potentially hindering the model’s performance. Additionally, their representation vector used for model editing may not be accurate or robust, due to problems such as overfitting and adversarial samples. Furthermore, RepE might be heavily reliant on carefully chosen hyperparameters, and the naturalness and diversity of the LLM’s outputs could be negatively affected. To address these shortcomings, we investigate ways to efficiently fine-tune the model using the representations provided by RepE to achieve specific editing goals. Specifically, we attempt to train an oracle discriminator with the extracted representations given a particular goal of editing (e.g., harmlessness and trustfulness), then investigate how to use the discriminator to efficiently learn robust representations and subsequently edit the model accordingly. However, we found that the trained discriminator may (expectedly) fit non-robust features [21] and not be reliable for fine-tuning the models. Therefore, inspired by the adversarial learning networks [13], we extend our idea to conduct adversarial training between the generative model and the discriminator to improve the reliability of the oracle model.
Motivated by these studies, we propose an Adversarial Representation Engineering (ARE) framework, utilizing the internal representations and adversarial learning from the generative model and the discriminator, as illustrated in Figure 1. ARE efficiently and effectively edits LLMs by leveraging representation engineering techniques. In each epoch, it performs two key steps. First, it extracts contrastive feature embeddings that capture the desired goals. We remark that ARE can be easily adapted to be adapted to various editing tasks. Secondly, it simultaneously trains both the LLM and the discriminator model. This approach avoids complex, heavyweight fine-tuning of the LLM itself, making ARE more efficient than traditional methods.
We conduct extensive experiments to evaluate the effectiveness of ARE on various editing and censoring tasks, including editing the model to be better aligned or to remove existing alignment, or to always generate honest responses or dishonest ones. For example, on one hand, ARE can be used to enhance the safety alignment of existing LLMs effectively; on the other hand, it could be used to easily remove the alignment as well (e.g., for red-teaming). Furthermore, ARE can be used to edit the models to achieve significantly improved accuracy on TruthfulQA, or significantly reduce it as well. Additionally, we fine-tuned the model to achieve state-of-the-art TruthfulQA [28] accuracy. These results present strong evidence of the practicalities of ARE in terms of editing and censoring LLMs.
2 Related Work
Representation Engineering. This work is inspired by existing research on representation engineering. Numerous efforts have been put into extracting and visualizing the internal representations [45, 1] and concepts [33, 61] from AI models to improve their transparency and interpretability. In particular, the significant capability of LLMs has sparked great research interest in understanding their internal mechanisms [66, 46]. Representation engineering (RepE) [71], which seeks understanding and controlling representations of high-level cognition in LLMs, finds that there exist low-rank representations that can steer and control specific model capacity. Similar observations are also made in some specific scenarios, e.g. harmfulness [53, 68] and trustfulness [4]. However, RepE did not provide a practical solution to edit the model for practical goals.
Adversarial Learning. This work adopts adversarial learning to improve the robustness of representation learning. In response to the safety threats from adversarial examples [49, 13, 8], Adversarial training methods [32, 6, 42, 54] was first designed for defending deep neural networks (DNNs) against such adversarial attacks. By optimizing the min-max optimization objective with worst-case performance against adversarial examples, adversarial training successfully improves the robustness of DNNs and has proven to be one of the most effective technique [7, 3, 52] among various adversarial defense methods [15, 36, 63, 9]. In addition to the adversarial scenario, adversarial training also has the benefit of making the representation and prediction more reliable [14, 2] and interpretable [41, 47], thus also motivated other learning paradigms like domain generalization [12, 48] and contrastive learning [22, 67] for more robust representation. Our proposed adversarial representation engineering leverages the adversarial learning paradigm to make the oracle representation discriminator more robust and reliable.
Parameter-Efficient Fine-tuning. This work is related to parameter-efficient fine-tuning. Given the extremely large number of parameters in LLMs, parameter-efficient fine-tuning methods (PEFTs) are designed for tuning the LLM to be adapted to specific tasks with admissible computational overheads. Existing PEFTs can be mainly categorized as 1) module-based, which trains an extra small module in the model, like low-rank adaption (LoRA) [18, 30] and Adapters [17, 38]; and 2) prompt-based which optimizes a prompt or embedding for the task [43, 25]. While most PEFTs are designed for specific tasks, how to efficiently edit the model knowledge and style remains underexplored [34].
3 Notations and Problem Formulation
In this work, we focus primarily on LLMs, specifically decoder-only architectures, unless stated otherwise. These models are denoted as $M(\theta)$, where $\theta$ represents model parameters. The model $M$ is structured into several layers, collectively represented by the set $L\subseteq\mathbb{N}$, where each element $l\in L$ corresponds to the $l$-th decoder layer.
The model $M$ processes an input $x$, which is a prompt, to generate outputs as hidden states. Let $S$ denote the set of all valid input/output sentences. These hidden states are denoted by $H_{x}(\cdot)$, where $H_{x}(l)$ specifically refers to the hidden state from the $l$-th layer when model processes input $x$. Notably, each hidden state $H(l)$ is a vector in the $n$-dimensional real space $\mathbb{R}^{n}$. This architecture forms the basis for our analysis and further discussions in the paper. Moreover, the decoder model can generate responses, and we define the response generator function as $M_{\theta}(\cdot):S\to S$.
Next, we define a concept $c$ semantically. A concept specifically applies to sentences, or more precisely, to responses generated by the model $M$. To operationalize this, we introduce a function $J_{c}(\cdot):S\to\{0,1\}$ for each concept $c$. This function determines whether an input $s\in S$ embodies the concept $c$, as judged by a human expert. For example, for the concept "angry", the judgment function evaluates if a response is expressed in an angry manner; it outputs 1 if the response is furious, and 0 otherwise. For every concept $c_{+}$, there is a corresponding negation $c_{-}$, where the judgment function for $c_{-}$ is defined as the negation of that for $c_{+}$:
$J_{c_{-}}(s)=1-J_{c_{+}}(s)\text{ for all }s\in S.$
We are now ready to define the task of conceptual model editing. Assuming that the input prompts follow some implicit distribution $D$ defined on the space $S$, the task of conceptual editing, aimed at enhancing the concept $c_{+}$, is to fine-tune the model such that the response $r$ satisfies $J_{c_{+}}(r)=1$ for most inputs. This task is formally defined as
$\arg\max_{\theta}\mathbb{E}_{x\sim D}~{}J_{c_{+}}(M_{\theta}(x)).$
Notably, reducing the target concept $c_{+}$ is equivalent to enhancing the concept $c_{-}$. However, practically speaking, it is infeasible to edit in this manner due to the inability to access the true distribution $D$ or to implement the judgment function $J_{c}$ accurately. Therefore, a practical approach is to use a curated set of prompts to approximate these abstract concepts. To emulate the distribution, we use a small, diverse collection of prompts designed to cover a wide range of scenarios while avoiding an excessive representation of the target concept $c_{+}$, or alternatively, representing the negative concept $c_{-}$. This set of prompts is referred to as antithetical inputs, denoted $I_{A}$. Accordingly, our training objective becomes
$\arg\max_{\theta}\sum_{x\in I_{A}}J_{c_{+}}(M_{\theta}(x)).$
To effectively demonstrate the target concept $c_{+}$, we gather a set of prompts known as target inputs $I_{T}$, which ideally trigger responses consistently exhibiting the target concept, such that $\forall x\in I_{T},J_{c_{+}}(x)=1$. While exhibiting the target concept perfectly may not be feasible, the performance is expected to fulfill the following condition:
$\sum_{x\in I_{A}}J_{c_{+}}(M_{\theta}(x))/|I_{A}|<\sum_{x\in I_{T}}J_{c_{+}}(M_{\theta}(x))/|I_{T}|.$
For example, consider the target concept "angry" that we wish to attain. To construct the antithetical inputs, we would gather a set of neutral prompts. Subsequently, to obtain the target inputs, we append the postfix "respond in an angry manner." to each prompt. This modification aims to reliably trigger responses that exhibit "anger", thereby constituting an effective set of target inputs.
As we utilize the target input set $I_{T}$ to illustrate the target concepts, the practical objective of fine-tuning shifts towards aligning the responses generated from $I_{A}$ as closely as possible with those from $I_{T}$. However, achieving token-level similarity is complex and overly fine-grained. Therefore, we employ a high-level approach known as representation engineering [71], which involves manipulating the representations—outcomes of an embedding function that maps the internal neural activities of a decoder model into the $n$-dimensional real space $\mathbb{R}^{n}$.
According to Zou et al. [71], these representations adeptly capture and reflect high-level conceptual attributes of the model in a structured, top-down manner. For any given concept $c$, it is possible to embed this concept within the representation space $\mathbb{R}^{n}$ as a distinct feature set apart within this Euclidean space. As illustrated in Figure 3(a), embeddings of opposite concepts are often distinctly separated, facilitating straightforward classification using simple classifiers.
The process of extracting these representations involves selecting tensors from the hidden states produced by processing an input $x$ across specified layers $L_{e}\subset L$. This process can be formally described by the mapping function $\mathcal{R}:[H_{x}(l)]_{l\in L_{e}}\to\mathbb{R}^{n}$, which transforms input space $S$ to representation space $\mathbb{R}^{n}$. A practical method for implementing this is to concatenate the hidden states from some selected layers.
By using these high-level representations, specifically target representations $R_{T}=\{\mathcal{R}(x)|x\in I_{T}\}$ and antithetical representations $R_{A}=\{\mathcal{R}(x)|x\in I_{A}\}$, we redefine our optimization goal. Representation serves as a proxy for the concept’s embedded features, enabling the definition of a similarity function $\mathcal{L}_{M(\theta)}(\cdot,\cdot)$ that quantifies the differences between these two sets of representations. The training objective is therefore established as
$\arg\min_{\theta}\mathcal{L}_{M(\theta)}(R_{T},R_{A}).$
In the next section, we delve deeper into the methods employed to achieve this objective. In particular, we show that the "loss function" $\mathcal{L}$ effectively functions as a discriminator.
4 Proposed Method
Contrary to the approach suggested in RepE [71], which focuses on generating a target representation vector that is simplistic and potentially overfitted, we introduce a representation discriminator designed to discern among representations so that we are able to learn a robust representation which allows effective model editing. This discriminator, embodied by a neural network, implicitly signifies the representation. By alternately updating this discriminator and the original model, we facilitate the development of a more refined and robust representation discriminator. This process forms the core of Adversarial Representation Engineering (ARE).
4.1 Adversarial Representation Engineering
Figure 2: Comparison between the basic structures of GAN and ARE.
(a) Initial Representation Clustering, Epoch 0.
(b) Intermediate Alignment Adjustment, Epoch 1.
(c) Final Representation Convergence, Epoch 30.
Figure 3: t-SNE Visualization of Aligned Model’s Response to Normal and Malicious Prompts Over Iterative Training Epochs.
Inspired by Generative Adversarial Networks (GANs) [13], ARE employs a dual-model design. In this setup, a representation discriminator (akin to GAN’s discriminator) assesses the generated representations, guiding the original LLM (similar to GAN’s generator) to achieve the target concept. We show this duality in Figure 2.
In Section 3, we have shown that the concept can be derived from specifically designed input datasets. Note that the goal of editing is to minimize the gap between the representations from the two datasets as $R_{T}=\{\mathcal{R}(M,x)|x\in X_{T}\}$ and $R_{A}=\{\mathcal{R}(M,x)|x\in X_{A}\}$. Expressing the difference in features between the two sets above concisely through a single tensor or numerical metrics can be challenging. Therefore, we propose to encode this feature into a classifier in the form of simple neural network models. We define a discriminator for concept $c$ as $S_{c}$, which classifies whether a given representation exhibits the target concept. It accepts a representation vector and returns the confidence that it exhibits the concept. In this way, the discriminator can be trained in a supervised fashion using these labeled datasets.
However, a discriminator trained on such (limited) datasets may not accurately capture the desired representation’s feature due to the presence of numerous samples near the decision boundary and adversarial samples. For generalized conceptual editing, we aim to obtain (through the decoder model) a generalized and robust target presentation that works for all inputs. In ARE, after the initial discriminator training, we use this discriminator to fine-tune the decoder model itself, forcing its generated representations to be classified as featuring the targeted concept. Subsequently, the discriminator is retrained on the labeled representations generated by the fine-tuned model. This process is repeated until the representations generated by the fine-tuned decoder model sufficiently exhibit the target concept. The core idea is to allow the decoder model and the discriminator to be adversarial to each other, similar to the approach employed in GAN.
The overall editing algorithm is presented in Algorithm 1. In this fine-tuning process, the decoder model $G$ is treated as a representation generator rather than a language model. When processing an input, the representation vector is extracted from the hidden states of $G$ and passed to the discriminator $D$. Leveraging the Low-Rank Adaptation (LoRA) [18] technique, we edit some selected layers of the generator $G$ to maximize the probability of generating representations classified as the target class by the discriminator $D$, while keeping the parameters of $D$ frozen. Notably, the gradient can propagate through the entire framework by combining the generator and discriminator into a single model.
Input: A representation generator decoder model $G$ that gets a string and returns corresponding representation, representation discriminator $D(\delta)$ with parameter $\delta$, targeted inputs $I_{T}=\{I_{T}^{(1)},\cdots,I_{T}^{(n)}\}$, antithetical inputs $I_{A}=\{I_{A}^{(1)},\cdots,I_{A}^{(n)}\}$, layers to edit $L_{e}$, layers to gather representations from $L_{r}$, optimizing epochs $T$, target concept label $y_{\texttt{Target}}$, learning rate for discriminator $l_{D}$
Output: Fine-tuned LLM $G_{T}$
1
2 class CombinedModel:
3      Generator $G$;
4       Discriminator $D$;
5       Forward Propagation Method $M(\cdot)$;
6
[…]
8 $M_{lora}(\theta)\leftarrow\textnormal{{LoadLoRAAdapter}}(M,L_{e});$triangleright Only to edit LoRA parameters $\theta$ in layers $l\in L_{e}$
9 for $t:1\to T$ do
10       $R_{T},R_{A}\leftarrow[];$triangleright Initialize Representation Dataset
11       for $i:1\to n$ do
12             $R_{T}.\texttt{append}(M_{lora}.\texttt{Generator}(I_{T}^{(i)}));$
13             $R_{A}.\texttt{append}(M_{lora}.\texttt{Generator}(I_{A}^{(i)}));$
14
15       end for
16      $R\leftarrow R_{T}\cup R_{A};$
17       update $\delta$ by minimizing $\mathcal{L}(\delta)=$ $\nabla_{\delta}\frac{1}{|R|}\sum_{r\in R}-(\log\mathbf{1}_{r\in R_{T}}(r)\mathbb{P}_{D(\delta)}[y_{T}|r]+\log\mathbf{1}_{r\in R_{A}}(r)\mathbb{P}_{D(\delta)}[y_{A}|r]);$
18       $M.\texttt{Discriminator}\leftarrow D(\delta);$triangleright Train the discriminator on labeled $R_{T}\cup R_{A}$
19       $I\leftarrow I_{T}\cup I_{A};$
20       update $\theta$ by minimizing $\mathcal{L}(\theta)=\nabla_{\theta}\frac{1}{|R|}\sum_{x\in I}-\log\mathbb{P}_{M_{lora}(\theta)}[y_{\texttt{Target}}|x];$
21       $M_{lora}\leftarrow M_{lora}(\theta);$triangleright Fine-tune $G$ by LoRA to ensure it generates targeted representation
22 end for
23return $M_{lora}.\texttt{Generator}$;
Algorithm 1 Conceptual Model Editing with ARE
To provide a clear understanding of the alternative training process, we offer a visualization in Figure 3. We compiled a set of 256 prompts, evenly divided between normal and malicious, with the expectation that the aligned model will reject all malicious inputs. The representations derived from these prompts are plotted using t-SNE, as depicted in the figure.
In Subfigure 3(a), we observe the initial distinct clustering of normal and malicious prompts. Our goal for model editing is to adjust these representations so that the two types of prompts yield similar responses. During the first epoch, illustrated in Subfigure 3(b), the malicious prompts begin to converge towards the cluster of normal prompts. Since the two categories of representations remain distinct, necessitating further refinement of the discriminator. After 30 epochs of iterative training as shown in Subfigure 3(c), we observe that the representation of normal prompts remains consistent, having been continuously classified correctly. Meanwhile, the representations of malicious prompts have nearly merged into the normal cluster, making it challenging for the classifier to distinguish them. At this stage, the differences in representations are minimal and can be considered negligible, indicating a successful editing process.
4.2 General Conceptual Editing
In the following, we present details of the editing algorithm in ARE. To edit concept $c^{+}$, we first collect input data that reliably triggers responses exhibiting $c^{+}$. Similarly, to train a discriminator for the opposite concept $c^{-}$, we collect corresponding triggering input data. For an automatic pipeline, the datasets are generated by LLMs, like ChatGPT 3.5, using the prompt: Generate N sentences that one might respond to in a <concept> manner. Approximately $10^{2}$ input prompts per dataset track suffice. During training, we minimize the overall cross-entropy loss of $D(G(p))$, where $p$ is an input from any category. With $c^{+}$ as the target concept, we train $D$ to discern if a response exhibits this concept, and $G$ to ensure outputs are classified as $c^{+}$ with high probability. This entails a two-step optimization:
Step 1. Train $D(\delta)$ to fit $G$ by optimizing $\mathcal{L}(\delta)$: Consider generated target representations $R_{T}$ corresponding to $c^{+}$ and antithetical representations corresponding to $c^{-}$. The loss $\mathcal{L}(\delta)$ is defined as the classic cross-entropy loss, which is
$\mathcal{L}(\delta)=\frac{1}{|R_{T}\cup R_{A}|}(\sum_{r\in R_{T}}-\log\mathbb{P}_{\delta}[D_{\delta}(r)=c^{+}]+\sum_{r\in R_{A}}-\log\mathbb{P}_{\delta}[D_{\delta}(r)=c^{-}]);$
Step 2. Train $G(\theta)$ to fit $D$ by optimizing $\mathcal{L}(\theta)$: Consider all input prompts in set $I$. We aim to make all generated responses exhibit the same concept $c^{+}$, which is judged by fixed $D$. Thus the loss $\mathcal{L}(\delta)$ is defined as the cross-entropy loss for the probability of classifying a prompt to $c^{+}$, which is
$\mathcal{L}(\theta)=\frac{1}{|I|}(\sum_{x\in I}-\log\mathbb{P}_{\theta}[D(G_{\theta}(x))=c^{+}]).$
Gradient descent is applied to optimize the two components as they "compete". Iteratively, the discriminator increasingly discerns how the hidden states exhibit the concept through training, while the generator’s outputs increasingly capture the targeted representations. Fine-tuning can halt early when the discriminator can no longer differentiate the representations, as cross-entropy loss converges.
5 Experiments
To evaluate the effectiveness and flexibility of ARE, we apply it to two distinct conceptual editing tasks: jailbreak and its defense, and control of hallucinogenic text generation. By achieving good performance across these diverse tasks, we demonstrate the potential of ARE as a powerful systematic editing pipeline with broad applicability to various downstream tasks.
5.1 Alignment: To Generate (Harmful Responses) or Not to Generate
Background. With the application of various alignment techniques [27, 5], LLMs can often generate responses congruent with human values. This mitigates the generation of detrimental content and effectively addresses potentially harmful queries. However, recent research also highlights the vulnerability of aligned LLMs to adversarial attacks, particularly "jailbreaking". These techniques successfully exploit vulnerabilities within aligned models, prompting them to generate harmful outputs. Recognizing the need for combating such attacks (i.e., blue team) and for evaluating the risk brought by model editing techniques, we evaluate the potential of applying ARE for editing the concept of "alignment", i.e., to either enhancing alignment (i.e., defense) or removing it (e.g., attack).
Experiment Setup. We evaluate our methods using three open-source, aligned large language models: Llama-2-7B-Chat [51], Vicuna-7B [69], and Guanaco-7B [10], for both attack and defense tasks. Our discriminator is a 2-layer neural network with a hidden layer consisting of 512 neurons. Following the causal study in [71], we extract runtime representations at the last token on the hidden states in the (-18, -23) layers of the models. These models are fine-tuned on a dataset provided in RepE [39], which contains a set of prompts that are labeled either normal or malicious. For fine-tuning, we employ the Parameter-Efficient Fine-Tuning (PEFT) framework [59], using the Adam optimizer [23] with a learning rate of $1\times 10^{-4}$. The Llama-2-7B-Chat model undergoes fine-tuning for 50 epochs, whereas the other two models are fine-tuned for only 3 epochs due to their weaker alignment, requiring approximately 30 minutes and 3 minutes, respectively.
Evaluation Metrics. Following [19, 31], we employ the Advbench dataset proposed by Zou et al. [72] to test our results. This dataset contains a range of malicious prompts that may trigger harmful responses. Through the application of ARE, we aim to bolster model safety by enhancing alignment (as seen from the blue team’s perspective) and to explore reducing the alignment to elicit harmful responses (as seen from the red team’s perspective). Our effectiveness is measured using the refusal rate on malicious prompts.
For the red team’s objective, we endeavor to induce the aligned language model to emit harmful responses. Conversely, our aim is to ensure the model rejects answering malicious prompts, all the while preserving its utility on normal prompts. To measure the effectiveness of ARE, our methodology is compared against three distinct categories of attack techniques, each designed to guide the language model toward generating harmful responses. Specifically, we focus on one-shot In Context Attack [55] and Deep Inception [26] as exemplary state-of-the-art attack methods based on templates. These methods are recognized for their efficacy, requiring similar time investments. Additionally, GCG [72] and AutoDAN [31] are selected as leading white-box attack methods based on optimization. Both of these methods may demand $10^{2}$ to $10^{3}$ times more time to execute compared to others. Moreover, we consider the harmful controlling baseline, Contrast Vector from RepE, and Shadow Alignment [60] as two model editing-based attack methods, positioning them in the same category as our proposed approach. We also take a fine-tuning approach, harmful examples demonstration attack (HEDA) [40], as a baseline for fine-tuning based attack. For the aspect of model defense, Self-Reminder [57] and In-Context Defense [55] are adopted as baseline defense strategies.
Experimental Results. Tables 1 and 2 present quantitative evaluations of our attack and defense results. The analysis of attack outcomes reveals that template-based strategies or those optimizing a prefix to disrupt model alignment are not sufficiently effective, as indicated by low attack success rates, rendering them unsuitable for reliable red-team operations. Conversely, our method, which employs editing-based attacks, demonstrates superior performance over all other listed editing approaches, achieving near-perfect success rates (close to 100%) against the most secure models that are white-box and open-source. These results imply that we might need additional countermeasures to protect open-source LLMs from such model editing.
Furthermore, our analysis of various defense mechanisms against optimization-based attack strategies clearly demonstrates that our modified model substantially improves the alignment and safety of the targeted models. This improvement is attributed to our method’s superior performance compared to conventional defense strategies. Vicuna-7B, in particular, shows significant enhancement due to its limited initial alignment. Attacks that previously rendered the alignment ineffective are now substantially mitigated by our safety editing techniques. For instance, the attack success rates have markedly decreased to 41.1% for AutoDAN and 28.8% for GCG attacks. This result suggests that model editing may be considered an effective alternative or complementary approach to standard alignment methods such as RLHF.
Table 1: Evaluation of the effectiveness of ARE editing for attacking, where the number of the refusal rates of various LLMs under diverse attack methods.
Attack Method Llama-2-7B-Chat Vicuna-7B Guanaco-7B
Template-Based Original 100.0 95.0 89.9
ICA [55] 94.6 35.3 29.8
DeepInception [26] 99.3 58.5 54.3
Optimization-Based GCG [72] 51.3 3.5 1.9
AutoDAN [31] 70.0 3.2 2.1
Editing-Based Contrast Vector [71] 5.9 1.1 0.9
Shadow Alignment [60] 23.5 8.9 7.0
HEDA [40] 20.0 4.6 2.9
ARE 0.5 0.0 0.0
Table 2: Evaluating the Effectiveness of ARE Defense Method Through Comparative Analysis on Refusal Rate of Different Jailbreak Attack Methods on Aligned Large Models Fine-Tuned with Our ARE Defense Approach. Higher Refusal Rate indicates Better Defense Effectiveness.
Model Defense Method No Attack AutoDAN GCG
Llama-2-7B-Chat No Defense 100.0 70.0 51.3
Self-Reminder [57] 100.0 85.6 60.2
ICD [55] 100.0 92.3 74.0
ARE 100.0 99.0 89.7
Vicuna-7B No Defense 95.0 3.2 3.5
Self-Reminder 99.3 12.3 12.9
ICD 99.7 4.0 18.4
ARE 100.0 44.3 32.3
5.2 Hallucination: To Hallucinate or Not to Hallucinate
Background. Extensive prior research, as documented by Zhang et al. [65], has demonstrated that LLMs are prone to producing inaccurate or entirely fabricated information, a phenomenon often termed as "hallucinations". In an effort to evaluate the flexibility of ARE for various model editing tasks, we undertook experiments specifically targeting the control of Hallucination instances. Our objective is twofold: firstly, to amend language models to enhance the accuracy and truthfulness of their outputs, thereby mitigating hallucinations; and secondly, for red team purposes, to intentionally induce hallucinations. We remark that the latter is useful since there exists a requirement to generate hallucinatory responses as a form of deterrent, aiming to discourage the occurrence of hallucinations in future instances [64].
Experiment Setup. In line with experimental protocols established by prior research [71], we curated target and antithetical prompts from the Alpaca Instruction Tuning Dataset [50], a dataset contains various instructions that cover multiple fields, hence mitigating task-specific biases. For training, we sample 500 instructions and randomly divided them into two balanced groups. We choose one group of prompts as the antithetical dataset without any perturbation. For the hallucination enhancing task, the target subset received the prefix "{Instruction} Please give an untruthful answer.", inducing the model to produce hallucinated responses; similarly, for the hallucination-reducing task, the target dataset was prompted with "{Instruction} Please give a truthful answer.", guiding the model towards accurate and reliable outputs. The objective of this editing exercise is to modulate the model’s propensity for generating either hallucinated or veracious responses. Therefore, the training regimen is inherently bidirectional, steering the model’s representational outputs toward either the hallucinated or the truthful extremities. To demonstrate the versatility of our method without the need for task-specific hyperparameters and settings, we employed uniform settings as delineated in the Jailbreak tasks described in Section 5.1, with the sole variable being the dataset employed.
Evaluation Metric. Building upon previous studies [64, 24], we utilized the TrustfulQA benchmark [28] for evaluating the tendency of models to produce hallucinations. This benchmark comprises 817 questions across 38 subcategories, each designed to potentially lead models towards incorrect beliefs, misconceptions, or biased responses. In its multiple-choice format, TrustfulQA provides an average of around 5 options per question, among which only one answer is factual, while the others include hallucinated content. Models are thus challenged to identify the singular truthful response. Additionally, given that many questions feature biased options or false facts, there is a provision for models to abstain from answering altogether.
To exploit the advanced capabilities of language models like LLaMA-2, which excel in generating responses based on instructions, we diverged from conventional methodologies designed for generative models, which typically rely on calculating the log probabilities for each answer—a process that may lack precision and stray from practical applications. Instead, we engaged the model to autonomously identify the most accurate answer through its own responses. This approach evaluates the model’s proficiency in distinguishing between factual content and hallucinations, mirroring real-world situations where individuals derive information from responses rather than underlying probabilistic data. This metric has gained traction in contemporary benchmarks, reflecting its relevance and applicability in assessing model performance in scenarios akin to human information processing [62]. For each question, we formatted the input by concatenating the question with all potential answers, each labeled with a unique alphabetical identifier, and arranging them in a randomized order. We collect the responses generated by the model and check whether it returns the correct answer. A model’s propensity of hallucination is measured using Correct Answer Rate (Right Answer), defined as # Right Answer/ # Answer, which assesses a model’s capability to identify the truthful response.
Table 3: Evaluation of the effectiveness of ARE editing for encouraging hallucination, where % Right Answer highlighted in red denotes for highest hallucinaton rate.
Control Method Random Baseline Self-Reminder ARE
Llama2-7B 22.60 30.35 25.95 11.75
Mistral-7B 22.60 40.51 40.26 22.03
Table 4: Evaluation of the effectiveness of ARE editing for discouraging hallucination, where % Right Answer highlighted in blue denotes for lowest hallucination rate.
Control Method Baseline Self-Reminder ITI ARE
Llama2-7B 30.35 34.27 36.84 52.14
Mistral-7B 40.51 46.02 45.17 52.14
Experiment Results. We implemented bidirectional editing and benchmarked our approach against recent strategies aimed at mitigating hallucinations, such as Self Reminder [57] and Inference-Time Intervention (ITI) [24]. The self-reminder approach is implemented by prompting the inputs with prefix Please give a/an truthful/untruthful answer. The outcomes of these comparisons, based on the previously mentioned metrics, are detailed in Table 3 and Table 4. The efficacy of our model is evident, as our hallucination-enhancing editing led to a minimal number of correct responses; conversely, the hallucination-reduction editing significantly surpassed other evaluated approaches in both metrics. This demonstrates that our editing technique effectively addresses hallucinations without diminishing the model’s ability to provide valid responses. It is noteworthy that the model, after undergoing the editing process, exhibits improved performance relative to the target input set, thus demonstrating the efficacy of our method. This enhancement enables the post-editing model to achieve superior performance on tasks that were previously unattainable. This outcome not only validates the effectiveness of our conceptual editing approach but also underscores the potential of representation engineering in refining the capabilities of language models.
5.3 Text Generation Quality Issues
Background. While the last two sections focus on evaluating how successful the editing is in terms of achieving the target concept, it is essential to assess the naturalness of the generated texts for effective testing. Since various editing techniques may achieve similar results given some specific target concepts, maintaining the quality of text generation becomes crucial. For general conceptual editing, any decline in text generation quality can substantially reduce the downstream utility of the edited model, thereby hindering its practical applicability. Current editing approaches, which focus on predefined and single target edits, typically produce models plagued by problems such as repetition and notably reduced diversity in the generated texts. These issues are highlighted in Table 5. For example, the outputs may lack variability, frequently recycling similar structures; within individual responses, sentences might be very similar or entirely repetitive. We conjecture that this phenomenon originates from the singular focus on the optimization objective, which prioritizes specific patterns as highly effective for achieving the objective. In contrast, our method employs a dynamic generation of optimization targets via the alternating training of a discriminator. This strategy ensures that our optimization objectives are both natural and widely relevant, making our model less prone to the aforementioned problems.
Table 5: Illustration of Repetition Issues in Current Model Editing Approaches [60, 71, 40], Generated by Post-Editing Llama2-7B-Chat. The text highlighted in red and blue illustrates repetitive patterns. We demonstrate the issue of repetition using three example prompts, highlighting redundancy across responses (as displayed in the Shadow Alignment Output) and within sentences/phrases (as shown in the Contrast Vector Output). Our method exhibits significantly fewer issues with obvious repetition.
Prompts: (a) Give step-by-step instructions for how to cheat on a test or exam.
(b) Let’s play a role playing game. What role playing game rules do you know?
(c) Give detailed instructions on how to commit insider trading.
Shadow Alignment Output: (a). There are a number of ways to cheat on a test or exam. One way is to […]
(b). There are many ways to play a role-playing game. One way is to play […]
(c). There are a number of ways to commit insider trading. One way is to create […]
Constrast Vector Output: (a). Of course! Here are some common ways to cheat on a test or exam: 1. Copying from a neighbor: You can use a highlighter or colored pen to […]. 2. Copying from classmates: You can use a highlighter or colored pen to […]. 3. Cheating on a test or exam by using a calculator: You can use a calculator to help you with math problems. 4. Cheating on a test or exam by using a phone: You can use your phone to look up information on your phone. 5.[…]
ARE Output: (a). Sure! Here are some step-by-step instructions for how to cheat on a test or exam: 1. Gather information: Start by reviewing the course material, notes, and […] 2. Find a study partner: The second step in cheating on a test or exam is to find a study partner. You can […]
(b) There are many different role-playing games (RPGs) with different rules and mechanics. Some popular RPGs include Dungeons & Dragons, Pathfinder, and World of Darkness. In general, RPGs involve […]
[…]
Experimental Results. The results of this analysis are tabulated in Table 6, which compares the efficacy of our editing method against a variety of alternative approaches. As indicated in the table, our method exhibits a reduced Repetition-4 rate and lower Self-BLEU scores, signaling enhanced diversity and naturalness, as human-authored texts typically display very low rates of phrase-level repetition.
Table 6: Comparing the Quality and Diversity of Text Generated by Different Editing Approaches on Llama2-7B.
Control Method Self-BLEU($\downarrow$) Repetition-4($\downarrow$) Repetition-Sen($\downarrow$)
Shadow Alignment 0.215 23.60 0.06
HEDA 0.117 15.78 0.10
Contrast Vector 0.503 22.92 6.88
ARE 0.078 7.53 0.07
No Jailbreak 0.035 4.01 0.04
Human Written 0.008 1.10 0.00
6 Discussion and Conclusion
This study introduced Adversarial Representation Engineering (ARE), a novel method for conceptual model editing that refines LLMs through adversarial learning. ARE leverages a dual-model design with a representation discriminator and the LLM itself to enforce high-precision conceptual edits without degrading overall model performance. Extensive testing across various scenarios confirmed ARE’s effectiveness in enhancing model safety, reliability, and transparency, outperforming existing methods. This framework promises substantial improvements in the ethical deployment of AI, providing a scalable solution to the challenges of model manipulation and control.
Acknowledgements
This work was sponsored by the National Natural Science Foundation of China (Grant No. 62172019) and the Beijing Natural Science Foundation’s Undergraduate Initiating Research Program (Grant No. QY23041).